OpenStudy (ali2x2):

answer these questions while i go brb 1 + 1 2 + 2 3 + 3 and so on

Elsa213 (elsa213):

1+1=11 2+2=22 3+3=33

OpenStudy (ali2x2):

:o

OpenStudy (ali2x2):

thanks i got an f :D

iYuko (iyuko):

math GENIUS

OpenStudy (ali2x2):

IKR

OpenStudy (ali2x2):

brb

Elsa213 (elsa213):

:D CONGRATULATIONS